For 589-nm light, calculate the critical angle for the following materials surrounded by air.

(a)cubic zirconia(n=2.20)
Answer #1

a) θ=Sin-1(1/2.2)=27 degrees

b)θ=Sin-1(1/1.66)=37 degrees

c)θ=Sin-1(1/1.309)=50 degrees
Answer #2

The critical angle is the angle of incidence above which total internal reflection occurs. The angle of incidence is measured with respect to the normal at the refractive boundary.

a) θ=Sin-1(1/2.2)=27 o

b)θ=Sin-1(1/1.66)=37 o

c)θ=Sin-1(1/1.309)=49.8 o